OpenStudy (anonymous):

Factor by grouping. b^3-2+5ab^3-10a

Directrix (directrix):

b^3-2+5ab^3-10a = 5ab^3 + b^3 - 10a - 2 = b^3 ( 5a + 1) - 2(5a +1) = ? Note that the common factor is (5a + 1)